OpenStudy (anonymous):

PLEASE HELP MEDAL/FAN?! How can I create a logarithmic function? What logarithmic function represents the data in the table? x f(x) 216 3 1296 4 7776 5

OpenStudy (anonymous):

@thomaster @undeadknight26 can you help?

undeadknight26 (undeadknight26):

didn't learn this yet sorry.

OpenStudy (anonymous):

its ok :D thanks for coming anyway

undeadknight26 (undeadknight26):

wells now let me try some people :) @amistre64 @divu.mkr @skullpatrol @shadowfiend @robtobey @texaschic101

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@surjithayer do u know this?

OpenStudy (anonymous):

\[216=6^3\] \[\log_{6} 216=3\] \[1296=6^4\] \[\log_{6} 1296=4\] \[7776=6^5\] \[\log_{6} 7776=5\]

OpenStudy (anonymous):

so \[f(x) = \log _{6}x\] right?

OpenStudy (anonymous):

thank you :D